# Generative Refocusing: Flexible Defocus Control from a Single Image

### v2

@@ 5 Conclusion
We present Generative Refocusing, a two-stage diffusion framework turning single images into controllable virtual cameras. By decoupling into DeblurNet (deblurring) and BokehNet (bokeh synthesis), our approach handles arbitrary inputs with intuitive controls over focus plane, bokeh intensity, and aperture shape. Through semi-supervised training on synthetic and real unpaired bokeh data, GenRefocus learns authentic optical behavior beyond simulators. Experiments on DPDD, RealDOF, LF-Bokeh, and LF-Refocus show consistent improvements, enabling text-guided refinement and aperture-shape editing.
Limitations and future work.
GenRefocus relies on monocular depth estimation; severe failures result in mislocalized focal planes. Complex aperture shapes require simulator-driven training data. Future work should generalize to richer aperture vocabularies, including user-drawn designs.
Appendix A Appendix Overview
In this appendix, we provide additional details on the training datasets used in our framework, the construction pipeline of our PointLight-1K dataset, a comparison with the VLM model, and visual comparisons from the BokehNet training ablation study. We also present failure cases and additional in-the-wild results on the accompanying HTML page.
Figure 9: Comparison with VLM model. Qualitative refocusing results comparing our Generative Refocusing framework with Gemini 3 Nano Banana Pro [20] given prompt ”focus on the man on the right”.
Appendix B Visual Comparison for BokehNet Training Ablation
In the main paper, we compare different training strategies for BokehNet. As illustrated in Fig. 10, under a foreground-bokeh setting with comparable bokeh strength in both rows, the model in (a), which is trained only on synthetic pairs, inherits artifacts from the simulator [48] and fails to correctly blur fine structures such as hair along object boundaries. In contrast, the model in (b), which is additionally trained with unpaired real bokeh data, produces smoother and more realistic blur around these edges.
Appendix C Training Datasets
We summarize the datasets used to train DeblurNet and BokehNet.
C.1 DeblurNet
For training DeblurNet, we use the DPDD [1] dataset and adopt all images from its official training split. This dataset provides paired defocused and all-in-focus images, which we use to supervise the deblurring stage.
We further leverage the RealBokeh_3MP [60] dataset, which contains images captured with relatively small apertures and thus does not strictly provide all-in-focus views. To obtain images that are sufficiently sharp for our purposes, we compute the Laplacian variance of each image as a simple focus measure and rank the dataset accordingly. We then retain the top $3000$ images with the highest Laplacian variance as additional training data for DeblurNet.
C.2 BokehNet
For training BokehNet, we require diverse all-in-focus inputs. We therefore draw all-in-focus images from the [82] and [30] collections. As in the DeblurNet stage, we compute the Laplacian variance for each candidate image and filter out blurry examples. After this filtering, we obtain a pool of approximately 1.71.7k all-in-focus images for simulating bokeh images. The handling of the real bokeh datasets BokehDiffusion [19] and RealBokeh_3MP [60] for bokeh synthesis is already described in the main paper.
Appendix D PointLight-1K Collection Pipeline
@@
Appendix E Comparison with VLM model
@@
We also compare our method with the recently released vision–language model Gemini-3 Nano Banana Pro [20], which explicitly supports text-driven refocusing. For this model, we use the prompt “focus on the man on the right” to specify the desired focus point. As shown in Fig. 9, Gemini-3 Nano Banana Pro can change the focus to some extent, but it also alters the facial expressions and appearances of people in both the foreground and the background. In contrast, our method preserves the original content while producing a realistic bokeh effect and accurate refocusing.
Figure 10: Visual Comparison for BokehNet Training Ablation.
Appendix F Failure Cases
We observe two main types of unsatisfactory refocusing results: (1) incorrect hallucinated details and (2) inaccurate bokeh blur. For the first type, even though DeblurNet can generate visually plausible high-frequency details, in cases of severe input blur it may hallucinate incorrect detail (see Fig. 11). For the second type, the bokeh blur can be imperfect because our defocus map is derived from monocular depth estimator [7]. When the estimated depth scale is biased, the resulting defocus map may fail to fully cover the intended focus region or may assign too much or too little blur to certain areas. Although our semi-supervised training strategy for BokehNet mitigates these issues to some extent, residual discrepancies with respect to the ground truth can still be observed in challenging cases.
Figure 11: Failure case. Due to the severely blurred input, the model hallucinates an incorrect time, changing 11:30 a.m. to 12:30 a.m.

### v3

@@ 5 Conclusion
We present Generative Refocusing, a two-stage diffusion framework turning single images into controllable virtual cameras. By decoupling into DeblurNet (deblurring) and BokehNet (bokeh synthesis), our approach handles arbitrary inputs with intuitive controls over focus plane, bokeh intensity, and aperture shape. Through training on various data sources, GenRefocus learns authentic optical behavior beyond simulators. Experiments on all benchmarks show consistent improvements, enabling aperture-shape editing.
5.0.1 Limitations and future work.
GenRefocus relies on the defocus map; Imperfections in the depth estimate can propagate to the defocus map and yield incorrect blur assignment. Transparent surfaces are particularly challenging, as monocular depth typically captures the depth of the front transparent surface while ignoring the background seen through it (see Fig. 9). In cases of severe blur, our diffusion-based method may fail to recover fine details and instead hallucinate textures (see Fig. 9), rendering it unsuitable for precision-critical fields such as scientific imaging or safety-sensitive applications. Complex aperture shapes require simulator-driven training data. Future work should generalize to richer aperture vocabularies, including user-drawn designs.
Appendix 0.A Overview
In the supplementary material, we provide additional details regarding the training datasets used in our framework and the construction pipeline of PointLight-1K dataset. We also present extended comparisons with DiffCamera [69], Bokehlicious [57], and Vision-Language Models. Furthermore, we provide qualitative results demonstrating the controllability of our approach, along with further analysis on ITW dataset [19] and a demonstration of the pre-deblur module’s effectiveness. We also present additional in-the-wild results on the accompanying HTML page.
Appendix 0.B Additional Training Details
In this section, we detail how we process datasets for training DeblurNet and BokehNet.
0.B.1 DeblurNet Training Data
DeblurNet is trained using a combination of paired defocus-blur data and high-quality sharp images. Specifically, we adopt all images from the official training split of the DPDD [1] dataset, which provides supervised pairs of defocused and all-in-focus images. To supplement this, we incorporate data from RealBokeh_3MP [57]. Since this dataset contains images with varying degrees of focus, we apply a quality filter by computing the Laplacian variance of each image as a focus measure. By ranking the dataset based on this metric, we retain the top 3,000 sharpest images to serve as additional supervision for the deblurring task.
0.B.2 BokehNet Training Data
Training BokehNet requires diverse, high-resolution all-in-focus inputs to simulate realistic bokeh effects. We draw candidate images from [80] and the EBB [27] collections, and similar to the DeblurNet stage, we use Laplacian variance to filter out blurry examples. This yields a refined pool of approximately 1.7K sharp images. Furthermore, we utilize a comprehensive dataset of 26K real bokeh images. This collection comprises 13K previously filtered and verified images from the ITW dataset [19], alongside 13K images newly curated for this work. The newly collected data consists of focus-consistent series captured with varying apertures, containing 2 to 4 images per set. To establish accurate ground truth for these series, we first conducted a manual verification and refinement process on the in-focus masks. This annotation step required 4 to 8 seconds per image, amounting to approximately 8 hours of manual effort in total. Utilizing these masks, we then optimized the parameter $K$ using simulator [43]. After obtaining the optimized $K$ values, we applied a Structural Similarity (SSIM) threshold to filter out sub-optimal results, ensuring that only reliable pairs are used for supervision.
Appendix 0.C PointLight-1K Collection Pipeline
@@
Table 7: Quantitative comparison on the DPDD dataset using $512\times 512$ central crops for the defocus deblurring task. Both methods are evaluated without any tiling strategy to ensure a strictly fair comparison of core architectural capacity. The best results are highlighted in bold.
Method LPIPS $\downarrow$ DISTS $\downarrow$ CLIP-IQA $\uparrow$ MANIQA $\uparrow$ MUSIQ $\uparrow$
DiffCamera [69] 0.2833 0.1793 0.4602 0.3502 43.2997
Ours 0.1285 0.1045 0.4637 0.3765 47.2070
Appendix 0.D Additional Comparison with DiffCamera
To ensure a strictly fair comparison of core model performance, we evaluate both methods under identical input conditions. Because DiffCamera [69] is constrained to its native $512\times 512$ training resolution and a fixed aspect ratio, we extract $512\times 512$ central crops from the DPDD benchmark. This guarantees that both models are evaluated at DiffCamera’s optimal setting.
As shown in Table 7, our method consistently outperforms DiffCamera across all full-reference and no-reference metrics. Notably, our approach achieves a significant reduction in LPIPS (0.1285 vs. 0.2833) and DISTS (0.1045 vs. 0.1793). This comprehensive superiority under a strictly controlled baseline confirms our robust generative capacity and highlights the critical advantage of leveraging real-world data during training.
Figure 10: Comparison with Bokehlicious [57]. Results are shown for historical images (top three rows) and in-the-wild images (bottom three rows). Bokehlicious [57] either fails to remove the blur entirely or introduces severe structural artifacts. Conversely, our proposed method exhibits strong generalization capabilities, robustly handling diverse blur conditions to synthesize crisp and visually pleasing outputs.
Appendix 0.E Qualitative comparison with Bokehlicious
In the main paper, we state that our model demonstrates greater robustness to diverse input conditions. To further substantiate this claim, we provide additional qualitative comparisons with Bokehlicious [57] in Fig. 10.
A primary real-world application of defocus deblurring lies in restoring invaluable legacy photographs that can no longer be recaptured, as well as rescuing modern in-the-wild images suffering from accidental misfocus. As shown in the first three rows of Fig. 10, we evaluate the models on historical images, including monochrome photographs. Under these conditions, Bokehlicious either yields negligible deblurring effects or introduces severe structural artifacts. Similarly, in the in-the-wild scenarios (the last three rows), Bokehlicious fails to effectively recover the sharp latent images and suffers from visual degradation.
In contrast, our proposed method successfully restores details across all these highly diverse conditions. From aging black-and-white photos to everyday digital captures, our model consistently produces visually appealing results, thereby validating its strong generalization capabilities and practical applicability.
Figure 11: Qualitative comparison with VLM model. Qualitative refocusing results comparing our Generative Refocusing framework with Gemini 3 Nano Banana Pro [20] given prompt "focus on the man on the right".
Appendix 0.F Qualitative comparison with VLM model
@@
We also compare our method with the vision–language model Nano Banana Pro [20], which explicitly supports text-driven refocusing. For this model, we use the prompt “focus on the man on the right” to specify the desired focus point. As shown in Fig. 11, Nano Banana Pro can change the focus to some extent, but it also alters the facial expressions and appearances of people in both the foreground and the background. In contrast, our method preserves the original content while producing a realistic bokeh effect and accurate refocusing.
Figure 12: Qualitative evaluation of controllability. Refocusing results across different blur levels ($K\in\{0,5,10,15\}$). The blue boxes highlight specific text regions to illustrate the defocus effect. As $K$ increases, the blurriness of the scene increases monotonically, demonstrating our model’s precise control over the bokeh effect.
Appendix 0.G Qualitative Results for Controllability
In the main paper, we quantitatively demonstrated that our framework outperforms all other diffusion-based baselines in terms of controllability. In this section, we qualitatively evaluate this controllability by examining the influence of the blur parameter $K$ and the defocus map. As shown in figure 12, the blurriness of the scene increases monotonically as the given $K$ becomes larger. This monotonic transition effectively proves that our model achieves precise controllability.
Figure 13: Camera model distribution. Percentage of images captured by the top 30 camera models in our dataset. Less frequent models are aggregated into the “Others” category.
Figure 14: Focal length distribution. The percentage distribution of focal lengths across ITW dataset.
Figure 15: Aperture (F-number) distribution. The percentage distribution of aperture settings across ITW dataset. Labels are presented as vertical fractions ($\frac{f}{N}$) for clarity.
Figure 16: Pixel ratio distribution. The percentage distribution of the pixel-to-sensor ratio across ITW dataset. This ratio, defined as the image’s largest edge length divided by the physical sensor width ($\text{px}/\text{mm}$).
Appendix 0.H ITW Dataset Analysis
This dataset is utilized to estimate the image bokeh level $K$ via EXIF metadata. We provide further analysis regarding the dataset’s diversity. Crucially, all data consists of real optical bokeh images captured directly by digital cameras, rather than artificial defocus synthesized through computational aperture techniques. In addition to the hardware diversity represented by the various camera models (Fig. 13), we analyze the distributions of the key parameters used in our $K$ estimation. Specifically, we provide the distributions for focal length (Fig. 14), F-number (Fig. 15), and pixel ratio (Fig. 16).
Figure 17: The effect of the pre-deblur module. Even when the pre-deblurred image is still distorted and blurry, using it as a condition yields higher-fidelity results (e.g., recovering “NEW YORK”) than removing this condition, which leads to distorted characters.
Appendix 0.I The effect of the pre-deblur module.
We introduce a variant that conditions DeblurNet on a pre-deblurred estimate. This provides a structural prior to disambiguate severe blur and improve fidelity. As shown in Fig. 17, even if the pre-deblurred image remains distorted, it helps the model recover fine details (e.g., the text “NEW YORK”).